# Pushing the Limits of Sparsity: A Bag of Tricks for Extreme Pruning

### v2

@@ 4.3 Ablation studies
4.3.1 Isolating DyReLU and Weight Sharing.
We first augment the standard RigL pipeline with only DyReLU phasing, replacing standard ReLU with DyReLU at initialization and gradually transitioning back to ReLU before the first learning rate schedule update. Weight sharing and cyclic sparsity are excluded in this isolated experiment. The results, summarized in Table 3 demonstrate a clear performance boost when utilizing DyReLU phasing, particularly at higher sparsity levels. At 99.99%, RigL completely loses its expressive capacity. However, as soon as DyReLU Phasing is used as a drop-in replacement for ReLU, it immediately starts converging. We observe that not only is training improved while DyReLU is in effect, but the performance continues to improve even after DyReLU has been completely phased out.
@@
ResNet34 RigL w/ ReLU 81.47 10.03
@@
ResNet50 RigL w/ ReLU 30.14 10.00
@@
Figure 5: Ablation study using CIFAR-100 and ResNet34. Impact of cyclic sparsity on density (left) and validation accuracy (right) compared to static sparsity. All runs are performed using the same settings except cyclic sparsity schedule.
@@
4.3.2 Effect of Cyclic Sparsity
Cyclic sparsity provides a flexible framework that allows networks to explore a wider parameter space. We perform an ablation study on the effect of cyclic sparsities using ResNet34 and CIFAR100 with a target sparsity of 99.95%. As shown in Figure 5, different cyclic schedules - whether starting from high sparsity and decreasing, or from low sparsity and increasing, consistently outperform the static sparsity across all configurations. While each schedule can be tuned differently (e.g., altitude, length of cycle, etc.), we consistently observe improved parameter exploration and generalization.
We conclude that having some form of cyclic sparsity improves performance over static sparsity. However, the large number of possible configurations make systematic experimentation challenging. Future work will focus on tuning these schedules to identify optimal schedule for different datasets and architectures.
@@ 5 Conclusion
In this paper, we have presented EAST, a new DST approach with a strong focus on extreme sparsities that incorporates DyReLU phasing, weight sharing, and cyclic sparsity scheduling. Through extensive empirical evaluation on CIFAR-10, CIFAR-100 and ImageNet using ResNet-34 and ResNet-50, we demonstrated EAST’s ability to achieve competitive performance at extreme sparsities, surpassing existing methods. Each component of EAST individually contributes to maintaining robust gradient flow and optimizing parameter exploration at extreme sparsity, while their combined effect can further enhance performance stability. Our results highlight the potential in practical applicability in resource-constrained environments, establishing it as a scalable and effective method for highly compressed networks. This work paves the way for future research to consider and further optimize models at extreme sparsities, an area that previous studies have often overlooked or deemed infeasible.
Acknowledgments: Our work was supported by UK EPSRC (EP/V042270/1), a University PhD studentship, and HPC Maxwell.
@@ A.1 Dynamic Sparse Training
where $\|\mathcal{M}\|_{0}$ is the number of active (non-zero) weights in the mask $\mathcal{M}$ and $\|\theta\|_{0}$ is the total number of weights in a dense network. We follow the same criteria for updating connections used in RigL [4]. First, we prune by magnitude, where the smallest magnitude weights are pruned. When pruning, the top-K operation selects the weights with the smallest absolute values:
@@ A.2 Dynamic ReLU
B. Proof
Here we provide the proof for the theorem in Section 3.1, which states that during the DyReLU to ReLU transition, the gradient flow is greater than the standard ReLU. We consider the two scenarios - before DyReLU phasing and during DyReLU phasing,
Before DyReLU phasing ($\beta=1$), we can express the network as $f(x;\beta(0))=f_{\text{DyReLU}}(x)$. Given the definition of DyReLU in A.2, the gradient using DyReLU can be expressed as:
$\frac{\partial y_{c}}{\partial x}=\frac{\partial(a_{k}^{c}(x)x_{c}+b_{k}^{c}(x))}{\partial x}$ (1)
Applying product rule to $a_{k}^{c}(x)x_{c}$ and chain rule:
$a_{k}^{c}(x)x_{c}=a_{k}^{c}\frac{\partial x_{c}}{dx}+x_{c}\frac{\partial a_{k}^{c}}{\partial x}$ (2)
$\frac{\partial y_{c}}{\partial x}=a_{k}^{c}(x)\frac{\partial x_{c}}{\partial x}+x_{c}\frac{\partial a_{k}^{c}(x)}{\partial x}+\frac{\partial b_{k}^{c}(x)}{\partial x}$ (3)
In DyReLU-B, $a_{k}^{c}$ and $b_{k}^{c}$ are negligible, $\frac{\partial a_{k}^{c}(x)}{\partial x}=0$ and $\frac{\partial b_{k}^{c}(x)}{\partial(x)}=0$. Therefore:
$\frac{\partial y_{c}}{\partial x}=a_{k}^{c}(x)\frac{\partial x_{c}}{\partial x}=a_{k}^{c}(x)$ (4)
For comparison, we express the gradients of both ReLU and DyReLU. The ReLU gradient is:
$\frac{\partial\text{ReLU}(x_{c})}{\partial x_{c}}=\begin{cases}1&\text{if }x_{c}>0,\\ 0&\text{if }x_{c}\leq 0\end{cases}$ (5)
Case 1: $x>0$. We first establish the gradient norm equations for ReLU and DyReLU as follows:
$\|\nabla\text{ReLU}(x)\|=\sqrt{\sum_{x^{c}>0}1^{2}}$ (6)
$\|\nabla\text{DyRELU}(x)\|=\sqrt{\sum_{x^{c}>0}(a_{k}^{c})^{2}}$ (7)
Since $a_{k}^{c}$ can be greater, equal to or less than 1, we rely on statistical evidence more than half are positive with a slope greater than 1 or negative [1], as well as our observation from Figure 4. Therefore, in a significant number of cases $a_{k}^{c}>1$, and DyReLU provides a greater gradient than the standard ReLU. Since $\sqrt{\sum_{x^{c}>0}(a_{k}^{c})^{2}}>\sqrt{\sum_{x^{c}>0}1^{2}}$, we have $\|\nabla\text{DyRELU}(x)\|>\|\nabla\text{ReLU}(x)\|$.
Case 2: $x\leq 0$. While $\|\nabla\text{ReLU}(x)\|=0$, DyReLU provides non-zero gradients through non-zero slopes $a_{k}^{c}$. When computing the gradient norm following Equation 7, since slopes $a_{k}^{c}$ can be non-zero for negative inputs (as observed empirically in the paper), and $(a_{k}^{c})^{2}>0$ for any non-zero slope regardless of the sign, we have $\|\nabla\text{DyReLU}(x)\|>0=\|\nabla\text{ReLU}(x)\|$.
During DyReLU phasing ($\beta\in(0,1)$), we have:
$f(x;\beta(t))=\beta(t)f_{\text{DyReLU}}(x)+(1-\beta(t))f_{\text{ReLU}}(x)$ (8)
The gradient at any point during transition is:
$\|\nabla f(x;\beta)\|$ $=\|\beta\nabla\text{DyReLU}(x)+(1-\beta)\nabla\text{ReLU}(x)\|$ (9)
$=\beta\|\nabla\text{DyReLU}(x)\|+(1-\beta)\|\nabla\text{ReLU}(x)\|$
$=\beta(\|\nabla\text{DyReLU}(x)\|-\|\nabla\text{ReLU}(x)\|)+\|\nabla\text{ReLU}(x)\|$
$>\|\nabla\text{ReLU}(x)\|$
C. Implementation Details
We share the hyperparameters used from Section 4.1 and 4.2 in Table C.1.
@@ C. Implementation Details
Table C.1: Experiment hyperparameters used in this paper. Batch Size (BS), Learning Rate (LR), Epochs, Learning Rate Drop (LR Drop), Optimizer with Momentum(Opt), Weight Decay (WD), Sparse Initialization (Init), Cyclic Sparsity-Related hyperparameters - Number of Cycles (n), Length of Cycle (l), Maximum Parameter Multiplier (m), Update Frequency during Cyclic Sparsity ($\Delta T_{cs}$), and Update Frequency for regular DST ($\Delta T_{dst}$).

### v3

@@ 4.2 Ablation Studies
Isolating DyReLU and Weight Sharing. We first augment the standard RigL pipeline with only DyReLU phasing, replacing standard ReLU with DyReLU at initialization and gradually transitioning back to ReLU before the first learning rate schedule update. Weight sharing and cyclic sparsity are excluded in this isolated experiment. The results, summarized in Table 4 demonstrate a clear performance boost when utilizing DyReLU phasing, particularly at higher sparsity levels. At 99.99%, RigL completely loses its expressive capacity. However, as soon as DyReLU Phasing is used as a drop-in replacement for ReLU, it immediately starts converging. We observe that not only is training improved while DyReLU is in effect, but the performance continues to improve even after DyReLU has been completely phased out.
@@
ResNet-34 RigL w/ ReLU 81.47 10.03
@@
ResNet-50 RigL w/ ReLU 30.14 10.00
@@
Effect of Cyclic Sparsity. Cyclic sparsity provides a flexible framework that allows networks to explore a wider parameter space. We perform an ablation study on the effect of cyclic sparsities using ResNet-34 and CIFAR-100 with a target sparsity of 99.95%. As shown in Figure 6, different cyclic schedules - whether starting from high sparsity and decreasing, or from low sparsity and increasing, consistently outperform the static sparsity (red) across all configurations. While each schedule can be tuned differently (e.g., altitude, length of cycle, etc.), we consistently observe improved parameter exploration and generalization. We conclude that having some form of cyclic sparsity may improve the performance over static sparsity. However, the large number of possible configurations makes systematic experimentation challenging. Future work will focus on tuning these schedules to identify optimal schedules for different datasets and architectures.
Figure 6: Ablation study using CIFAR-100 and ResNet-34. Impact of cyclic sparsity on density (left) and validation accuracy (right) compared to static sparsity. All runs are performed using the same settings except cyclic sparsity schedule.
@@ 5 Conclusion
In this paper, we introduced EAST, a new DST approach focusing on extreme sparsities that incorporates DyReLU phasing, weight sharing, and cyclic sparsity scheduling. Through extensive empirical evaluation on CIFAR-10, CIFAR-100 and ImageNet using ResNet-34 and ResNet-50, we demonstrated EAST’s ability to achieve competitive performance at extreme sparsities, surpassing existing methods by a large margin. Each component of EAST individually contributes to maintaining robust gradient flow and optimizing parameter exploration at extreme sparsity, while their combined effect can further enhance performance stability. Our results highlight the potential for practical applicability in resource-constrained environments, establishing it as a scalable and effective method for highly compressed networks. This work opens up future research to optimize models at extreme sparsities, an area often overlooked in previous studies.
Acknowledgments. Our work was supported by the UKRI EPSRC funding council (EP/V042270/1), a University of Aberdeen PhD studentship, the HPC facility ”Maxwell”, and the Interdisciplinary Institute.
Supplementary Material
Section A: Preliminaries
@@ A.1 Dynamic Sparse Training
where $\|\mathcal{M}\|_{0}$ is the number of active (non-zero) weights in the mask $\mathcal{M}$ and $\|\theta\|_{0}$ is the total number of weights in a dense network. We follow the same criteria for updating connections used in RigL. First, we prune by magnitude, where the smallest magnitude weights are pruned. When pruning, the top-K operation selects the weights with the smallest absolute values:
@@ A.2 Dynamic ReLU
A.3 Phasing out DyReLU
$\beta(t)=1-\frac{t-t_{\text{start}}}{t_{\text{end}}-t_{\text{start}}}$
$f(x;t=0)=f_{DyReLU}(x)$
$f(x;t)=\beta(t)\cdot f_{DyReLU}(x)+(1-\beta(t))\cdot f_{ReLU}(x)$
By the end of phasing $t=T_{\text{end}}$, we have:
$f(x;t=T_{\text{end}})=f_{ReLU}(x)$
B. Implementation Details
We share the hyperparameters used from Section 4 in Table B.1.
Table B.1: Experiment hyperparameters used in this paper. Batch Size (BS), Learning Rate (LR), Epochs, Learning Rate Drop (LR Drop), Optimizer with Momentum(Opt), Weight Decay (WD), Sparse Initialization (Init), Cyclic Sparsity-Related hyperparameters - Number of Cycles (n), Length of Cycle (l), Maximum Parameter Multiplier (m), Update Frequency during Cyclic Sparsity ($\Delta T_{cs}$), and Update Frequency for regular DST ($\Delta T_{dst}$).